OpenStudy (anonymous):

Solve using the square root property: 4a^2-45+0

OpenStudy (anonymous):

-45?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

if this equals 0, then the answer is sqrt(45/4)

OpenStudy (anonymous):

=0*

OpenStudy (mathstudent55):

4a^2 - 45 = 0 4a^2 = 45 a^2 = 45/4 a = +-sqrt(45/4) a = +-3sqrt(5)/2